# RNAGenScape: Property-guided Optimization and Interpolation of mRNA Sequences with Manifold Langevin Dynamics

### v1

@@ Abstract
mRNA design and optimization are important in synthetic biology and therapeutic development, but remain understudied in machine learning. Systematic optimization of mRNAs is hindered by the scarce and imbalanced data as well as complex sequence-function relationships. We present RNAGenScape, a property-guided manifold Langevin dynamics framework that iteratively updates mRNA sequences within a learned latent manifold. RNAGenScape combines an organized autoencoder, which structures the latent space by target properties for efficient and biologically plausible exploration, with a manifold projector that contracts each step of update back to the manifold. RNAGenScape supports property-guided optimization and smooth interpolation between sequences, while remaining robust under scarce and undersampled data, and ensuring that intermediate products are close to the viable mRNA manifold. Across three real mRNA datasets, RNAGenScape improves the target properties with high success rates and efficiency, outperforming various generative or optimization methods developed for proteins or non-biological data. By providing continuous, data-aligned trajectories that reveal how edits influence function, RNAGenScape establishes a scalable paradigm for controllable mRNA design and latent space exploration in mRNA sequence modeling.
@@ 1 Introduction
Messenger ribonucleic acids (mRNAs) design and optimization are important (Qin et al., 2022; Metkar et al., 2024) but remain understudied in machine learning (Castillo-Hair & Seelig, 2021; Schlusser et al., 2024). Even small edits to mRNA sequences can strongly affect their stability, translation efficiency, and eventual protein output (Zhang et al., 2023; Li et al., 2025). For instance, modifying the non-coding 5’ untranslated region (UTR) can tune the degradation rate of transcripts and therefore regulate protein production (Castillo-Hair et al., 2024; Ma et al., 2024). Such guided controls are directly relevant to applications such as mRNA vaccines (Pardi et al., 2018; Chaudhary et al., 2021) and protein replacement therapies (Qin et al., 2022; Vavilis et al., 2023), where improving translation efficiency or stability can improve efficacy and reduce dosage. However, systematic optimization in the mRNA space remains an open challenge, because (1) viable mRNAs occupy only a narrow subset of the vast ambient design space (Zhang et al., 2023; Calvanese et al., 2024), (2) data collected in this field are scarce and imbalanced, with many regions sparse or undersampled (Taubert et al., 2023; Asim et al., 2025), and (3) the sequence-function relationships are highly complex (Licatalosi & Darnell, 2010; Weinreb et al., 2016).
We present RNAGenScape, a property-guided Manifold Langevin dynamics framework dedicated for mRNA sequence design and optimization. RNAGenScape consists of three key components. 1 An organized autoencoder (OAE) that learns the manifold of mRNA sequences and organizes the space by the target property, enabling efficient exploration within a biologically plausible subspace rather than the ambient sequence space. 2 A manifold projector that contracts each update back onto the learned manifold, preserving biological plausibility. By preprocessing the data with SUGAR (Lindenbaum et al., 2018) to augment undersampled regions by filling “holes” in the manifold, we ensure that the manifold projector remains effective even under sparse training conditions (e.g., as few as 2,000 data points in one dataset), addressing a common challenge in mRNA data. 3 A property-guided manifold Langevin dynamics that leverages OAE and manifold projector to optimize target properties of the mRNAs, smoothly interpolate between sequences, and ensure that all intermediate results remain close to the viable mRNA manifold.
Across three real-world mRNA datasets that span two orders of magnitude in size, RNAGenScape consistently improved target properties with high success rates and efficiency, outperforming existing approaches originally developed for proteins or generic sequence data.
In summary, our main contributions are as follows.
  1. 1.
    Framework: We propose RNAGenScape, a manifold Langevin dynamics framework that enables interpolation and continuous property-guided optimization of mRNA sequences starting from real data points, offering biologically grounded sequence modeling.
  2. 2.
    Manifold constraint: We introduce a learned manifold projector that ensures biological plausibility throughout optimization trajectories.
  3. 3.
    Efficiency: Unlike diffusion-based models which typically start from Gaussian noise and explore the entire Euclidean space, we restrict our exploration to the manifold and start from existing sequences, allowing faster training and inference.
  4. 4.
    Empirical validation: We provide results on three real mRNA datasets, demonstrating that our method improves target properties while maintaining manifold fidelity, outperforming various optimization and generation methods.
@@ 2 Preliminaries
2.1 Manifold hypothesis and manifold learning
The manifold hypothesis (Cayton et al., 2008; Narayanan & Mitter, 2010; Fefferman et al., 2016) posits that high-dimensional data lie near a low-dimensional manifold embedded in the ambient space. Formally, each observation $x_{i}\in\mathbb{R}^{n}$ arises from a smooth nonlinear map $\mathbf{f}:\mathcal{M}^{d}\to\mathbb{R}^{n}$ applied to a latent variable $z_{i}\in\mathcal{M}^{d}$, where $d\ll n$.
Manifold learning methods seek to recover this latent structure by constructing representations that preserve intrinsic geometry (Van Dijk et al., 2018; Moon et al., 2019; Burkhardt et al., 2021; Liu et al., 2024; Liao et al., 2024; Liu et al., 2025a; b; Sun et al., 2025; Tong et al., 2023). Diffusion geometry (Coifman & Lafon, 2006; Van Dijk et al., 2018; Lindenbaum et al., 2018) provides one such paradigm, where local similarities are defined via a kernel on the pairwise similarities, and a Markov transition probability matrix is obtained by row normalization. This diffusion process encodes the intrinsic geometry of the data.
A point is considered on-manifold if it lies within the range of the nonlinear map $\mathbf{f}$, while off-manifold points deviate from this structure and may correspond to invalid or adversarial samples (Rifai et al., 2011; Li et al., 2023). Thus, projecting updated points back to the manifold is critical for robustness and geometry-aware optimization (He et al., 2023b).
@@ 2.2 Langevin-dynamics and beyond
[…]
Langevin Dynamics (Song & Ermon, 2019) has been employed in generative models to sample from high-dimensional data distributions using only an estimate of the score function $\nabla_{x}\log p(x)$. In particular, it first trains a neural network $\bm{s_{\theta}}$ to approximate the score function of data injected with Gaussian noise. Sampling is then performed via annealed Langevin dynamics, given by $\tilde{x}_{t}=\tilde{x}_{t-1}+\frac{\eta_{i}}{2}\bm{s_{\theta}}(\tilde{x}_{t-1},\sigma_{i})+\sqrt{\eta_{i}}\bm{z_{t}}$. Here, $\bm{s_{\theta}}(\tilde{x}_{t-1},\sigma_{i})$ is the learned score function at noise level $\sigma_{i}$, and $\eta_{i}$ is the step size at that level. By gradually annealing from high to low noise, this procedure enables generation of high-quality samples without an explicit likelihood or energy model.
3 RNAGenScape
The key components of our framework are 1 OAE: an autoencoder module whose latent space is organized by the target property (Section 3.1, Figure 1a), 2 Manifold Projector: a module that brings the updated latent embeddings back to the learned data manifold during each step of optimization/interpolation (Section 3.2, Figure 1b & 1d), and 3 Property-guided manifold Langevin dynamics, a procedure that integrates the two aforementioned modules to enable property optimization and interpolation (Section 3.3 & 3.4, Figure 1e).
Once trained, these components allows RNAGenScape to optimize the target property of a given sequence (Section 4.3) and interpolate between existing sequences (Section 4.5).
Figure 1: Schematic of RNAGenScape. (a) We first train an organized latent space for mRNA sequences by jointly optimizing reconstruction and property prediction objectives. (b) We then train a manifold projector while the encoder’s weights are frozen. (c) For undersampled mRNA manifolds, we use SUGAR to learn key dimensions in the manifold and fill undersampled regions. (d) During optimization, the manifold projector brings off-manifold points back to the manifold. (e) We can use the encoder and the manifold projector to optimize the properties of given input mRNA sequences or interpolate between sequences. Notably, the intermediate products can also be decoded. Best viewed zoomed in.
3.1 Learning a latent space organized by property
We begin by training an organized autoencoder (OAE), where the latent space is implicitly structured via supervision from a property prediction task (Figure 1a). Similar to a vanilla autoencoder (Hinton & Salakhutdinov, 2006), the encoder $\mathcal{E}$ maps the input mRNA sequence $x$ to a latent representation $z$, which is decoded by $\mathcal{D}$ back to the sequence space. In addition to this standard architecture, a predictor $\mathcal{P}$ infers properties $\hat{y}$ from the embedding $z$. Formally, $z_{i}=\mathcal{E}(x_{i}),\,\,\hat{x}_{i}=\mathcal{D}(z_{i}),\,\,\hat{y}_{i}=\mathcal{P}(z_{i})$.
The latent space $\mathcal{Z}$ is thus shaped by jointly optimizing the reconstruction loss and the prediction loss (equation 1), encouraging it to learn sequence-relevant information while being organized by the target properties. $\lambda_{\text{Pred}}$ and $\lambda_{\text{Recon}}$ are hyperparameters that balance between organizing the property landscape and capturing sequence information. They are empirically set to 1 and 5 in all experiments. Here, $x_{i}\in\mathbb{R}^{V}$ is the ground truth one-hot encoding of the nucleotide at position $i$ in sequence $x$ with an mRNA vocabulary of size $V$.
[…]
3.2 Training a manifold projector
@@
mRNA datasets are typically scarce, undersampled, and biased toward specific experimental conditions, which makes it difficult to learn a faithful latent manifold directly from the train data. To address this, we adopt SUGAR (Lindenbaum et al., 2018), a diffusion geometry-based generative method that learns the geometry of the data and samples the manifold uniformly. This augmentation enriches the latent space with geometry-preserving samples, helping the model better approximate the underlying mRNA manifold even in sparse regions.
@@
Learning the Manifold Projection
Algorithm 1 Manifold Projector
@@
To keep the generated trajectories aligned with the latent data manifold, we introduce a manifold projector $\Psi$. As illustrated in Figure 1b and magnified in Figure 1d, $\Psi$ takes in a noisy optimized point $\tilde{z}$ and projects it back onto or near the manifold.
To train $\Psi$, we adopt a denoising objective that contracts noisy samples back towards the clean points on the latent manifold. Given a clean latent embedding $z$, we construct a short corruption chain $\tilde{z}^{(0)}=z,\qquad\tilde{z}^{(k)}\sim C(\tilde{\mathcal{Z}}|\tilde{\mathcal{Z}}^{(k-1)},\sigma_{k}),\quad k=1,\dots,K,$ where $C(\cdot,\sigma_{k})$ denotes Gaussian corruption with noise level $\sigma_{k}$. The projector is trained to reverse each step by predicting $\tilde{z}^{(k-1)}$ from $\tilde{z}^{(k)}$, yielding the objective in equation 2.
$\mathcal{L}_{\Psi}\;=\;\mathbb{E}_{z\sim p_{\text{data}}}\;\sum_{k=1}^{K}\;\mathbb{E}_{\tilde{z}^{(k)}\sim C(\cdot\mid\tilde{z}^{(k-1)},\sigma_{k})}\Big[\,\|\Psi(\tilde{z}^{(k)})-\tilde{z}^{(k-1)}\|_{2}^{2}\,\Big]$ (2)
When $K=1$, this reduces to the standard denoising autoencoder loss (Vincent et al., 2008). In practice, we keep $K$ small (e.g. $1$ to $3$) to capture local updates near the data manifold, rather than simulating long diffusion chains from Gaussian noise. As a result, our algorithm is fast during training and inference.
@@ 3.3 Property-guided manifold Langevin dynamics
Given a trained encoder $\mathcal{E}$, a property predictor $\mathcal{P}$, and a manifold projector $\Psi$, our Langevin-dynamics framework optimizes sequences for a target property. Starting from the latent embedding $z=\mathcal{E}(x)$ of a sequence $x$, we iteratively update it using a gradient-based drift term $\nabla_{z}f(z)$, inject Gaussian noise $\epsilon$, and apply a manifold projection $\Psi(\cdot)$ to ensure biological plausibility and interpretability. We define the update rule as described in equation 3.
$z_{t+1}$ $=\Psi\!\left(z_{t}+\text{d}z_{t}\right),\quad\text{d}z_{t}=\tfrac{\eta}{\tau}\nabla_{z}f(z_{t})+\sqrt{2\eta}\,\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,I)\vskip-4.0pt$ (3)
@@
3.4 Interpolating between sequences
Beyond property optimization, our framework also enables interpolation between existing mRNA sequences by guiding the latent embedding of one sequence toward that of another. Specifically, given a source sequence $x_{\text{source}}$ and a target sequence $x_{\text{target}}$, we first obtain their latent embeddings via the encoder: $z_{\text{source}}=\mathcal{E}(x_{\text{source}})$ and $z_{\text{target}}=\mathcal{E}(x_{\text{target}})$.
Starting from $z=z_{\text{source}}$, we run property-guided manifold Langevin dynamics with an additional force term that pulls the latent toward $z_{\text{target}}$. The interpolation force is defined as $f_{\text{interp}}(z,z_{\text{target}})=-\frac{z-z_{\text{target}}}{\|z-z_{\text{target}}\|_{2}}$, which provides a normalized directional bias toward the target point. Incorporating this into the Langevin update changes the drift term while all other components remain intact, as described in equation 4.
$\text{d}z_{t}=\frac{\eta}{\tau}f_{\text{interp}}(z_{t},z_{\text{target}})+\sqrt{2\eta}\,\epsilon_{t}$ (4)
This modification steers the latent trajectory smoothly toward the target, enabling interpretable interpolations between biological sequences.

### v2

# RNAGenScape: Property-Guided, Optimized Generation of mRNA Sequences with Manifold Langevin Dynamics

@@ Abstract
Generating property-optimized mRNA sequences is central to applications such as vaccine design and protein replacement therapy, but remains challenging due to limited data, complex sequence-function relationships, and the narrow space of biologically viable sequences. Generative methods that drift away from the data manifold can yield sequences that fail to fold, translate poorly, or are otherwise nonfunctional. We present RNAGenScape, a property-guided manifold Langevin dynamics framework for mRNA sequence generation that operates directly on a learned manifold of real data. By performing iterative local optimization constrained to this manifold, RNAGenScape preserves biological viability, accesses reliable guidance, and avoids excursions into nonfunctional regions of the ambient sequence space. The framework integrates three components: 1 an autoencoder jointly trained with a property predictor to learn a property-organized latent manifold, 2 a denoising autoencoder that projects updates back onto the manifold, and 3 a property-guided Langevin dynamics procedure that performs optimization along the manifold. Across three real-world mRNA datasets spanning two orders of magnitude in size, RNAGenScape increases median property gain by up to 148% and success rate by up to 30% while ensuring biological viability of generated sequences, and achieves competitive inference efficiency relative to existing generative approaches.
@@ 1 Introduction
Messenger ribonucleic acids (mRNAs) have emerged as a core modality for modern therapeutics, enabling applications ranging from mRNA vaccine design and protein replacement therapy to gene regulation and synthetic biology (Pardi et al., 2018; Chaudhary et al., 2021; Qin et al., 2022; Vavilis et al., 2023). In these settings, the nucleotide sequence of an mRNA directly determines properties such as stability, translational efficiency, and ultimately protein yield. Even small sequence edits, particularly in non-coding regions such as the 5’ untranslated region (UTR), can lead to substantial changes in degradation rates and ribosome loading (Castillo-Hair et al., 2024; Ma et al., 2024; Li et al., 2025). As a result, systematic optimization of the mRNA sequence has the potential to improve therapeutic efficacy, reduce dosage requirements, and expand the design space of viable mRNA-based interventions.
Despite its importance, mRNA sequence generation and optimization remains underexplored from a machine learning perspective (Castillo-Hair and Seelig, 2021; Schlusser et al., 2024). A key challenge is that viable mRNA sequences occupy a narrow and structured subset of the astronomically large ambient sequence space (Calvanese et al., 2024; Zhang et al., 2023). Random or unconstrained edits can lead to nonfunctional transcripts, while experimentally measured datasets are typically small, imbalanced, and sparsely cover the space of interest (Taubert et al., 2023; Asim et al., 2025). Moreover, sequence-function relationships in mRNA are highly nonlinear and context-dependent, making the optimization of mRNA sequences a challenging problem (Licatalosi and Darnell, 2010; Weinreb et al., 2016).
To address the challenges, we introduce RNAGenScape, a property-guided manifold Langevin dynamics framework for mRNA sequence generation built around three key components. 1 An autoencoder and a property predictor, together termed an organized autoencoder (OAE), that jointly learn a latent manifold of mRNA sequences while organizing this manifold according to the target property. This latent manifold represents the subspace of biologically viable mRNA sequences rather than the much larger Euclidean ambient space. 2 A manifold projector, implemented as a denoising autoencoder, that maps data near the learned manifold onto the manifold, preventing drift toward implausible regions. To ensure robustness under sparse and imbalanced data regimes common in mRNA studies, we augment training data using SUGAR (Lindenbaum et al., 2018), which fills holes in the manifold and stabilizes manifold projection even with as few as 2,000 sequences. 3 A property-guided manifold Langevin dynamics that iteratively refines sequences by combining gradient information with stochastic exploration, while ensuring that sequences generated at every optimization step remain close to the viable mRNA manifold.
Our choice to perform iterative optimization directly on the learned data manifold, rather than adopting diffusion-based generation from Gaussian noise, is motivated by several practical considerations. First, mRNA sequence optimization is inherently local, where improvements are typically made relative to an existing sequence. Second, property-guided Langevin dynamics relies on gradients from a predictor trained on real data, and therefore staying close to the data manifold helps obtain stable and meaningful guidance. In contrast, standard diffusion and flow-matching models frequently traverse off-manifold regions during generation, leading to unstable or misleading guidance from predictors. Third, constraining the optimization trajectory to the manifold encourages the generated sequences to remain within biologically viable regions of sequence space. Finally, avoiding repeated diffusion to and denoising from a noise distribution far away from the data manifold reduces computational overhead and enables faster inference.
We evaluate RNAGenScape on three real-world mRNA datasets under diverse experimental settings. Across these datasets, RNAGenScape improves median property gain by up to 148% over state-of-the-art methods and success rates by up to 30%, while preserving biological viability and demonstrating generalization in a held-out optimization study. In addition, RNAGenScape is highly efficient with a throughput 68% higher than the fastest biological sequence optimization method.
@@ 2 Preliminaries
2.1 Manifold assumption and manifold learning
The manifold assumption (Cayton and others, 2008; Narayanan and Mitter, 2010; Fefferman et al., 2016) is a fundamental principle in machine learning which hypothesizes that high-dimensional natural data lie near a low-dimensional manifold embedded in the ambient space (Li and He, 2025). Formally, each observation $x_{i}\in\mathbb{R}^{n}$ arises from a smooth nonlinear map $\mathbf{f}:\mathcal{M}^{d}\to\mathbb{R}^{n}$ applied to a latent variable $z_{i}\in\mathcal{M}^{d}$, where $d\ll n$.
Manifold learning methods seek to recover this latent structure by constructing representations that preserve intrinsic geometry (Van Dijk et al., 2018; Moon et al., 2019; Burkhardt et al., 2021; Liu et al., 2024; Liao et al., 2024; Liu et al., 2025a, b; Sun et al., 2025; Tong et al., 2023; Zhang et al., 2025). A point is considered on-manifold if it lies within the range of the nonlinear map $\mathbf{f}$, while off-manifold points deviate from this structure and may correspond to invalid or adversarial samples (Rifai et al., 2011; Li et al., 2023). Thus, projecting updated points back to the manifold is critical for robustness and geometry-aware optimization (He et al., 2023b).
SUGAR (Lindenbaum et al., 2018) is a diffusion geometry-based method that learns the intrinsic geometry of the data manifold and samples uniformly along the learned geometry. When observed data are limited, augmenting training sets with SUGAR enriches the learned manifold with geometry-preserving samples, particularly in sparse regions.
@@ 2.2 Langevin-dynamics and beyond
Langevin Dynamics (Song and Ermon, 2019) has been employed in generative models to sample from high-dimensional data distributions using only an estimate of the score function $\nabla_{x}\log p(x)$. In particular, it first trains a neural network $\bm{s_{\theta}}$ to approximate the score function of data injected with Gaussian noise. Sampling is then performed via annealed Langevin dynamics, given by $\tilde{x}_{t}=\tilde{x}_{t-1}+\frac{\eta_{i}}{2}\bm{s_{\theta}}(\tilde{x}_{t-1},\sigma_{i})+\sqrt{\eta_{i}}\bm{z_{t}}$. Here, $\bm{s_{\theta}}(\tilde{x}_{t-1},\sigma_{i})$ is the learned score function at noise level $\sigma_{i}$, and $\eta_{i}$ is the step size at that level. By gradually annealing from high to low noise, this procedure enables generation of high-quality samples without an explicit likelihood or energy model.
Figure 1: Schematic of RNAGenScape. a, We encode a given input mRNA sequence into a latent manifold organized by property, and iteratively optimize the sequence along the property landscape. The manifold projector $\Psi$ ensures the optimization stays on or near the manifold. Notably, we can decode the intermediate optimization result after every step. b, The organized latent manifold is created by jointly optimizing reconstruction and property prediction objectives. c, The manifold projector is trained to project perturbed data points back to the manifold. d, For undersampled mRNA manifolds, we use SUGAR (Lindenbaum et al., 2018) to learn key dimensions in the manifold and fill undersampled regions. e, During optimization, the manifold projector brings off-manifold points back to the manifold.
3 RNAGenScape
The key components of our framework are 1 OAE: an autoencoder jointly trained with a prediction model to learn a latent manifold of mRNA sequences organized by the target property (Section 3.1 and Figure 1b), 2 Manifold Projector: a denoising autoencoder that brings the updated latent embeddings back to the learned data manifold during each step of optimization (Section 3.2 and Figure 1c, 1e), and 3 Property-guided manifold Langevin dynamics, a procedure that integrates the two aforementioned modules to enable property-guided generation (Section 3.3 and Figure 1a).
Once trained, these components allow RNAGenScape to optimize the target property, such as translation efficiency or stability, of a given mRNA sequence.
3.1 Learning a latent manifold organized by property
We begin by training an organized autoencoder (OAE), where the latent manifold is implicitly structured via supervision from a property prediction task (Figure 1b).
Similar to a vanilla autoencoder (Hinton and Salakhutdinov, 2006), the encoder $\mathcal{E}$ maps the input mRNA sequence $x$ to a latent representation $z$, which is decoded by $\mathcal{D}$ back to the sequence space. In addition to this standard architecture, a predictor $\mathcal{P}$ infers properties $\hat{y}$ from the embedding $z$. Formally, $z_{i}=\mathcal{E}(x_{i}),\,\,\hat{x}_{i}=\mathcal{D}(z_{i}),\,\,\hat{y}_{i}=\mathcal{P}(z_{i})$.
The latent manifold $\mathcal{Z}$ is thus shaped by jointly optimizing the reconstruction loss and the prediction loss (equation 1), encouraging it to learn sequence-relevant information while being organized by the target properties. $\lambda_{\text{Pred}}$ and $\lambda_{\text{Recon}}$ are hyperparameters that balance between organizing the property landscape and capturing sequence information. They are empirically set as described in Supplementary Section A. Here, $x_{i}\in\mathbb{R}^{V}$ is the ground truth one-hot encoding of the nucleotide at position $i$ in sequence $x$ with an mRNA vocabulary of size $V$.
@@ 3.1 Learning a latent manifold organized by property
3.2 Learning a module to project updates to manifold
@@
mRNA datasets are typically scarce and unevenly sampled, which can hinder learning a faithful latent manifold directly from observed data. To mitigate this issue, we augment the latent embeddings using SUGAR (Lindenbaum et al., 2018), a diffusion geometry-based method introduced in Section 2. This augmentation enriches the learned manifold with geometry-preserving samples, particularly in sparse regions, improving robustness of subsequent manifold projection.
@@
Training the manifold projector
Algorithm 1 Manifold Projector $\Psi$
@@
To access reliable property guidance during generation and keep the generated trajectories aligned with the latent data manifold, we introduce a manifold projector $\Psi$. As illustrated in Figure 1c and magnified in Figure 1e, $\Psi$ takes in a noisy optimized point $\tilde{z}$ and projects it back onto or near the manifold.
To train $\Psi$, we adopt a denoising objective that projects noisy samples back towards the clean points on the latent manifold. Given a clean latent embedding $z$, we construct a short corruption chain $\tilde{z}^{(0)}=z,\qquad\tilde{z}^{(k)}\sim C(\tilde{\mathcal{Z}}|\tilde{\mathcal{Z}}^{(k-1)},\sigma_{k}),\quad k=1,\dots,K,$ where $C(\cdot,\sigma_{k})$ denotes Gaussian corruption with noise level $\sigma_{k}$. The projector is trained to reverse each step by predicting $\tilde{z}^{(k-1)}$ from $\tilde{z}^{(k)}$, yielding the objective in equation 2.
$\mathcal{L}_{\Psi}=\mathbb{E}_{z\sim p_{\text{data}}}\sum_{k=1}^{K}\mathbb{E}_{\tilde{z}^{(k)}\sim C(\cdot\mid\tilde{z}^{(k-1)},\sigma_{k})}\Big[\,\|\Psi(\tilde{z}^{(k)})-\tilde{z}^{(k-1)}\|_{2}^{2}\,\Big]$
When $K=1$, this reduces to the standard denoising autoencoder loss (Vincent et al., 2008). We keep $K$ small (e.g. $1$ to $3$) to capture local updates near the data manifold, rather than simulating long diffusion chains from Gaussian noise. As a result, our algorithm is fast during training and inference.
@@ 3.3 Property-guided manifold Langevin dynamics
Given a trained encoder $\mathcal{E}$, a property predictor $\mathcal{P}$, and a manifold projector $\Psi$, our Langevin-dynamics framework optimizes sequences for a target property. Starting from the latent embedding $z=\mathcal{E}(x)$ of a sequence $x$, we iteratively update it using a gradient-based drift term $\nabla_{z}f(z)$, inject Gaussian noise $\epsilon$, and apply a manifold projector $\Psi(\cdot)$ to ensure biological viability and interpretability. We define the update rule as described in equation 3.
$\begin{split}z_{t+1}&=\Psi\!\left(z_{t}+\text{d}z_{t}\right)\\ \quad\text{d}z_{t}&=\tfrac{\eta}{\tau}\nabla_{z}f(z_{t})+\sqrt{2\eta}\,\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,I)\end{split}$ (3)